OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (katrinakaif):

Parallel means same slope therefore your 7 should be the same now substitute your given points into the equation (y - y1) = m (x-x1) (y +3) = 7 (x - 1)